Past records indicate that the probability of online retail orders that turn out to be fraudulent is 0.050.05. Suppose​ that, on a given​ day, 2020 online retail orders are placed. Assume that the number of online retail orders that turn out to be fraudulent is distributed as a binomial random variable. Complete parts​ (a) through​ (d) below. a. What are the mean and standard deviation of the number of online retail orders that turn out to be​ fraudulent?

. What are the mean and standard deviation of the number of online retail orders that turn out to be​ fraudulent?

. What is the probability that zero online retail orders will turn out to be​ fraudulent?

What is the probability that one online retail order will turn out to be​ fraudulent?

What is the probability that two or more online retail orders will turn out to be​ fraudulent?
Answer #1

a)

here mean of distribution=μ=np= 1
and standard deviation σ=sqrt(np(1-p))= 0.9747

b)  probability that zero online retail orders will turn out to be​ fraudulent=P(X=0)=20C0(0.05)0(0.95)20 =0.3585

c)

probability that one online retail order will turn out to be​ fraudulent=P(X=1)=20C2(0.05)1(0.95)19 =0.3774

d)

probability that two or more online retail orders will turn out to be​ fraudulent=P(X>=2)=1-P(X<=1)

=1-(P(X=0)+P(X=1)) =1-(0.3585+0.3774)=0.2642